### v1

# Deep-ELA: Deep Exploratory Landscape Analysis with Self-Supervised Pretrained Transformers for Single- and Multi-Objective Continuous Optimization Problems

Abstract
In many recent works, the potential of Exploratory Landscape Analysis (ELA) features to numerically characterize, in particular, single-objective continuous optimization problems has been demonstrated. These numerical features provide the input for all kinds of machine learning tasks on continuous optimization problems, ranging, i.a., from High-level Property Prediction to Automated Algorithm Selection and Automated Algorithm Configuration. Without ELA features, analyzing and understanding the characteristics of single-objective continuous optimization problems would be impossible.
Yet, despite their undisputed usefulness, ELA features suffer from several drawbacks. These include, in particular, (1.) a strong correlation between multiple features, as well as (2.) its very limited applicability to multi-objective continuous optimization problems. As a remedy, recent works proposed deep learning-based approaches as alternatives to ELA. In these works, e.g., point-cloud transformers were used to characterize an optimization problem’s fitness landscape. However, these approaches require a large amount of labeled training data.
Within this work, we propose a hybrid approach, Deep-ELA, which combines (the benefits of) deep learning and ELA features. Specifically, we pre-trained four transformers on millions of randomly generated optimization problems to learn deep representations of the landscapes of continuous single- and multi-objective optimization problems. Our proposed framework can either be used out-of-the-box for analyzing single- and multi-objective continuous optimization problems, or subsequently fine-tuned to various tasks focussing on algorithm behavior and problem understanding.
Keywords
Deep Learning, Exploratory Landscape Analysis, Single-Objective Optimization, Multi-Objective Optimization, Automated Algorithm Selection, High-level Property Prediction
1 Introduction and Related Work
Optimization problems, often found at the heart of numerous scientific and industrial applications, present challenges that necessitate robust and efficient problem-solving methodologies. A central concern in this field is the analysis and characterization of the continuous optimization problem landscapes. One methodological concept that has emerged as particularly significant for such a characterization is Exploratory Landscape Analysis (ELA; Mersmann et al.,, 2011). ELA facilitates a numerical representation of single-objective continuous optimization problems, providing insights into their intrinsic characteristics. The numerical features harvested from ELA serve as the foundation for various machine-learning tasks pertinent to continuous optimization. Such tasks include, but are not limited to, High-level Property Prediction (HL2P; Mersmann et al.,, 2011; Seiler et al.,, 2022; Volz et al.,, 2023), Automated Algorithm Selection (AAS; Rice,, 1976; Kerschke et al.,, 2019), and Automated Algorithm Configuration (AAC; Hutter et al.,, 2009; Huang et al.,, 2019; Schede et al.,, 2022).
While ELA has undoubtedly transformed how researchers approach single-objective continuous optimization problems, it is not without its challenges. Two of the most pronounced limitations include the strong correlations between numerous features and their restricted use in the multi-objective domain. The advent of deep learning has ushered in a myriad of solutions across disciplines. Recent studies have proposed leveraging deep learning, specifically point-cloud transformers, as a potential remedy to the shortcomings of ELA (Seiler et al.,, 2022; Prager et al.,, 2022). However, such approaches, while promising, come with their own set of challenges, most notably the requirement of vast amounts of labeled training data.
Given the strengths and weaknesses of both, ELA and deep learning-based methodologies, there is a compelling case to be made for a synthesis of the two. This paper seeks to bridge this gap by introducing a hybrid approach, capitalizing on the merits of both paradigms. We thus present a novel method that employs four pre-trained transformer models tailored to extract deep representations of the landscapes of both single- and multi-objective continuous optimization problems. This innovative approach promises enhanced flexibility and efficacy in the realm of optimization problem analysis and understanding.
We introduce these pre-trained transformers as Deep Exploratory Landscape Analysis (Deep-ELA). The strengths of Deep-ELA are prominently seen in its inherent adaptability to multi-objective continuous optimization problems and the reduced correlation among its learned features. We demonstrate the straightforward adaptability of Deep-ELA by applying it to rigorous tests across three distinct case studies that showcase their efficacy:
  1. 1.
    A classification task involving problem instances from the Black-box Optimization Benchmarking Suite (BBOB; Hansen et al.,, 2009). Here, we focus on modeling and predicting their High-Level Properties, effectively revisiting the experiments previously conducted by Mersmann et al., (2011) and Seiler et al., (2022).
  2. 2.
    An assessment of single-objective AAS on BBOB, featuring twelve algorithms sourced from the COmparing Continuous Optimisers (COCO; Hansen et al.,, 2019) platform. This particular study echoes the works of Kerschke and Trautmann, (2019) and later Prager et al., (2022).
  3. 3.
    An AAS study, applied to a set of multi-objective optimization problems, utilizing seven distinct algorithms. While this study draws inspiration from Rook et al., (2022), it ventures into a different direction, focusing on AAS rather than AAC.
The remainder of this paper is structured as follows. Initially, we equip readers with a concise background on ELA, providing an overview of feature-free alternatives and laying out the notation employed throughout the paper (Chapter 2). Subsequently, Chapter 3 dives into the architectural essence of Deep-ELA. Following this, we delve into a thorough exploration of the various datasets featured in this work in Chapter 4. We then pivot to a comparative analysis, contrasting the outcomes of Deep-ELA with traditional ELA across the previously mentioned case studies in Chapter 5. In Chapter 6, we present our conclusions and provide opportunities for future research.
2 Background
In the following, we will briefly outline the notation of this paper. An optimization function is given by:
$Z:\mathcal{X}\to\mathcal{Y},\quad\mathbf{x}\mapsto(z_{1}(\mathbf{x}),z_{2}(\mathbf{x}),\ldots,z_{m}(\mathbf{x}))^{\top}$ (1)
where $\mathcal{X}\subseteq\mathbb{R}^{d}$ is the decision and $\mathcal{Y}\subseteq\mathbb{R}^{m}$ the objective space. Every scalar function $z_{i}:\mathcal{X}\to\mathbb{R}$, with $i\in\{1,\ldots,m\}$, symbolizes a distinct objective function that maps from the decision space $\mathcal{X}$ to a real value. In this context, $\mathbf{x}\in\mathcal{X}$ is referred to as a candidate solution, and the set $X=\{\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{n}\}\subseteq\mathcal{X}$ is formed of $n$ such solutions. An objective or fitness vector is represented by $\mathbf{y}=Z(\mathbf{x})$ with $\mathbf{y}\in\mathcal{Y}$ and a length of $m$. If $m>1$, then $Z$ is a multi-objective function, whereas for $m=1$, $\mathbf{y}$ is a scalar, and $Z$ is a single-objective function. Finally, $Y=\{Z(\mathbf{x})\mid\mathbf{x}\in X\}$ is a set of objective values.
In single-objective optimization, the task is defined as:
$\mathbf{x}^{*}=\operatorname*{\text{arg\,min}}_{\mathbf{x}\in X}Z(\mathbf{x})$ (2)
Here, $\mathbf{x}^{*}$ is the optimal solution for $Z$. The goal of single-objective optimization is to find a solution $\mathbf{x}^{*}$ that w.l.o.g. minimizes the function $Z$ globally. If multiple optimal solutions exist, $Z$ is multimodal, and the aim shifts to locating all optimal solutions.
In contrast, in the multi-objective setting, optimizing for one objective can negatively affect another. Therefore, any solution provides a trade-off between the objectives, and the goal of multi-objective optimization is to find a set of optimal solutions, known as a non-dominated or Pareto set. Given two candidate solutions $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$, Pareto dominance, denoted as $\mathbf{x}_{i}\prec\mathbf{x}_{j}$, means $\mathbf{x}_{i}$ dominates $\mathbf{x}_{j}$, if it is equal or better in all objectives, and strictly better in at least one of them. Formally, this is represented as:
$z_{k}(\mathbf{x}_{i})$ $\leq z_{k}(\mathbf{x}_{j})\quad\forall k\in\{1,\ldots,m\},\text{ and }$ (3)
$z_{l}(\mathbf{x}_{i})$ $<z_{l}(\mathbf{x}_{j})\quad\exists\>l\in\{1,\ldots,m\}.$ (4)
The goal of multi-objective optimization is to determine the Pareto set
$X_{E}=\{\mathbf{x}_{i}\in X\mid\nexists\mathbf{x}_{j}\in X:\mathbf{x}_{j}\prec\mathbf{x}_{i}\}$ (5)
and its mapping to the objective space, called Pareto front
$Y_{E}=\{Z(\mathbf{x})\mid\mathbf{x}\in X_{E}\}.$ (6)
In many studies concerning algorithm behavior and problem understanding, $Z$ is treated as a blackbox optimization problem, implying that direct formal analyses cannot be conducted on $Z$. Instead, preliminary analyses are often carried out on a small sample of candidate solutions $X$, which is usually derived using a random or quasi-random generator like Latin Hypercube Sampling (LHS; McKay et al.,, 1979) or Sobol sequences (Sobol’,, 1967). The corresponding image $Y$ is obtained by evaluating the optimization problem $Z$ at the candidate solutions $X$. The tuple $(X,Y)$ is then available for subsequent studies.
Such studies span problem- and algorithm behavior-understanding, algorithm design, and AAS. Hereafter, we will refer to these types of studies as downstream tasks or downstream studies. This terminology is borrowed from the self-supervised training community, which differentiates between two phases: (1.) pre-training and (2.) fine-tuning. Self-supervised learning addresses tasks with insufficient training datasets. Models are thus pre-trained using an auxiliary training task – potentially using a synthetic set if the primary training set is inappropriate (e.g., too small or risk of overfitting). Subsequently, these models are applied or fine-tuned for the original downstream task.
2.1 Exploratory Landscape Analysis
Problem $I\in\mathcal{I}$ Comp. Features $\colon\mathcal{I}\rightarrow\mathcal{F}_{\text{ELA}}$ Select $\dot{S}\colon\mathcal{F}_{\text{ELA}}\rightarrow\mathcal{A}$ Algorithm Portfolio $\mathcal{A}$ Solve $I$ with $A$ (a) Feature-Based AAS
Problem $I\in\mathcal{I}$ Select $S\colon\mathcal{I}\rightarrow\mathcal{A}$ Algorithm Portfolio $\mathcal{A}$ Solve $I$ with $A$ (b) Feature-Free AAS
Figure 1: Comparison of the feature-based (left) versus feature-free (right) approach on one common downstream task: Automated Algorithm Selection (AAS). In the realm of AAS, there is no single universally superior algorithm for all problem instances. Instead, AAS uses a portfolio of algorithms $\mathcal{A}=\{A_{1},\ldots,A_{n_{\mathcal{A}}}\}$. The optimal selector is formally defined as $S\colon\mathcal{I}\rightarrow\mathcal{A}$. Typically, the selector $S$ is trained using machine learning to optimize a given performance metric. However, standard machine learners in AAS cannot process raw problem instances directly, necessitating a transformation into numerical vectors. This transformation is given as $F:\mathcal{I}\to\mathcal{F}\subseteq\mathbb{R}^{{n_{\mathcal{F}}}}$, where $F$ is a mapper converting an instance $I\in\mathcal{I}$ into a real-valued vector, termed (instance) features, in the feature space $\mathcal{F}$. Therefore, in a standard AAS scenario, the selector is defined as $\dot{S}\colon\mathcal{F}\rightarrow\mathcal{A}$, accepting features rather than actual instances.
For downstream tasks involving continuous optimization problems, traditional machine learning methods aim to learn a specific underlying task. In AAS, for instance, classical machine learners like Support Vector Machines (SVM; Cortes and Vapnik,, 1995), Random Forests (RF; Breiman,, 2001), or $k$-Nearest Neighbors ($k$NN; Cover and Hart,, 1967), are utilized to address the Algorithm Selection Problem (ASP) (see Figure 1a). However, these classical learners cannot directly process raw samples of candidate solutions, i.e., $(X,Y)$, as the learners necessitate feature vectors as input per problem instance. Yet, transforming the set of (randomly distributed) points into a vector is not suitable, as such a mapping either (a) lacks an unambiguous order of the points, or (b) requires the points to be aligned in a grid structure, which in turn is affected by the curse of dimensionality. Therefore, it is essential to convert the set of candidate solutions $(X,Y)$ into a vector of numerical values that provides a meaningful representation of the candidate tuple. This conversion, or transformation, is performed by a mapping function, or short mapper, $F:(\mathcal{X},\mathcal{Y})\to\mathcal{F}\subseteq\mathbb{R}^{{n_{\mathcal{F}}}}$, which translates the decision and objective values into a fixed-length real-valued vector, termed (instance) features, in the feature space $\mathcal{F}$.
For continuous optimization problems, Exploratory Landscape Analysis (ELA; Mersmann et al.,, 2011) features are commonly employed in many downstream tasks, including AAS. Formally, ELA is a function that maps samples $(X,Y)$ from the decision and objective space $(\mathcal{X},\mathcal{Y})$ to the ${c_{\text{ELA}}}$-dimensional feature space $\mathcal{F}_{\text{ELA}}\subseteq\mathbb{R}^{{c_{\text{ELA}}}}$
$\text{ELA}\colon(\mathcal{X},\mathcal{Y})\rightarrow\mathcal{F}_{\text{ELA}}.$
Over time, a plethora of ELA features for single-objective optimization problems emerged (Kerschke and Preuss,, 2023), such as:
Refer to caption (a) ELA Features
Refer to caption (b) Deep-ELA Features
Figure 2: Comparison of exemplary correlation matrices of (Deep-)ELA features on BBOB: (a) classical ELA features (here: meta-model, dispersion and $y$-distribution), and (b) Deep-ELA features (Large-$50d$). Correlations are calculated across all 24 functions of the BBOB suite, but individually for every instance $1$ to $20$ and dimension $2,3,5,10$. Afterward, the 80 correlation maps are mean-aggregated.
  1. 1.
    Classical ELA (Mersmann et al.,, 2011) is a superset of six different groups of ELA features, of which the meta-model- and y-distribution-features are the most commonly used ones.
  2. 2.
    Dispersion (Lunacek and Whitley,, 2006) measures the spread in decision space of the best fraction of points in relation to the spread of all points of the full sample.
  3. 3.
    Information Content (Muñoz et al.,, 2014) measures, i.a., the landscape’s smoothness and ruggedness based on statistics summarizing a series of random walks.
  4. 4.
    Fitness Distance Correlation (Jones et al.,, 1995) measures the correlation between distances in the objective and distances in the decision space.
  5. 5.
    Nearest Better Clustering (Kerschke et al.,, 2015) compares the relation between the set of nearest neighbors in the decision space to the set of nearest better neighbors, where ‘better’ relates to neighbors with lower objective values.
  6. 6.
    Miscellaneous (Kerschke and Trautmann,, 2019) is a collection of different feature types such as features based on Principal Component Analysis (PCA).
While the utility of ELA features is widely recognized, they also have some limitations (Kerschke and Preuss,, 2023). As, e.g., discussed by Renau et al., (2019) and Eftimov et al., (2020), many ELA features exhibit strong correlations to each other, as illustrated in Figure 2a, or have a low Signal to Noise Ratio (SNR), see Figure 3. Moreover, some features from sets like Classical ELA, Information Content, and PCA are sensitive to random scaling and shifting (Renau et al.,, 2020); see Prager and Trautmann, (2023) for an in-depth analysis of ELA feature sensitivity. These issues necessitate feature selection for many downstream tasks to eliminate noise or redundancy. Finally, ELA features are primarily tailored for single-objective optimization problems, not multi-objective ones. In fact, Kerschke and Trautmann, (2016) used ELA features to characterize bi-objective problems; however, they only considered simple test problems from DTLZ (Deb et al.,, 2005) and ZDT (Zitzler et al.,, 2000), and applied ELA to the single-objective components of these problems, neglecting any interaction between the objectives.
2.2 Deep Learning-Based Approaches as Alternative to ELA
Figure 3: Signal to Noise Ratio (SNR) of ELA features (left boxplot) on BBOB compared to the SNR values of features from four Deep-ELA models. We used $\text{SNR}=\mu^{2}/\sigma^{2}$ with mean $\mu$ and standard deviation $\sigma$. For $\sigma\simeq 0$, values are imputed with $10^{12}$, which is the highest observed value. Higher values indicate lower noise. SNR values are calculated per feature based on instances $1$ to $20$ and then mean-aggregated over the $24$ functions and four dimensions ($2,3,5,10$).
In the evolutionary computation (EC) community, methods that avert features are often termed feature-free. Yet, in the deep learning domain, this term can be misleading. In the latter, features concern internally learned features of complex deep learning models, whereas in the EC domain, (ELA) features refer to summary statistics quantifying the intrinsic characteristics of the given optimization problem (see Figure 1b for an illustration of feature-free AAS).
In recent studies, efforts have been made to solve downstream tasks without depending on ELA features, thus, sidestepping their limitations. In the context of discrete optimization (which is beyond the scope of this paper), Alissa et al., (2019) employed a Long Short-Term Memory (LSTM; Hochreiter and Schmidhuber,, 1997) for AAS on the $1d$-Bin Packing problem, without pre-computed features. Similarly, Seiler et al., (2020) used Convolutional Neural Networks (CNN; LeCun and Bengio,, 1995) for AAS on the Traveling Salesperson Problems (TSP) without pre-computed instance features.
Within continuous optimization, Prager et al., (2021) introduced a method applying CNNs to two-dimensional single-objective optimization problems. This was expanded by Seiler et al., (2022), who presented image- and point cloud-based techniques for high-dimensional, single-objective optimization problems. Notably, the Point Cloud Transformer (PCT; originally proposed by Guo et al.,, 2021) delivered top-tier results without the need for ELA features. Subsequently, Prager et al., (2022) employed PCT alongside traditional ELA-based methods for AAS on continuous, single-objective optimization tasks. Only ELA-based Multi-Layer Perceptrons (MLP) with a carefully designed loss function managed to surpass the performance of the PCTs.
Compared to deep learning-based feature-free techniques, ELA-based methods usually require significantly fewer training instances under the assumption that these smaller benchmark sets are nonetheless representative and comprehensive. This becomes particularly advantageous when optimization problems are scarce or crafting a training dataset is computationally expensive. While data augmentation can amplify the training set, deep learning methods may still generalize poorly in case of limited training data. Note that in this context, small means a few hundred and large implies tens or even hundreds of thousands of instances.
3 Deep Exploratory Landscape Analysis
In the following, we aim to merge the advantages of both feature-based and feature-free analysis in the context of continuous optimization. We introduce a large, pre-trained deep learning model trained on millions of randomly generated continuous, single- and multi-objective optimization instances. The generator for these tasks draws inspiration from Tian et al., (2020) and van Stein et al., (2023), but with adaptations to suit our requirements. Our model’s general goal is the automated extraction of instance features from initial samples of the given optimization instance. We dub this methodology Deep Exploratory Landscape Analysis (Deep-ELA).
Deep-ELA is trained to be invariant against scaling, shifts, and rotations, yielding higher SNR values and minimally correlated features (see Figure 2b and Figure 3) with high importance. Hence, feature selection becomes less important – though still potentially valuable in some cases – as will be demonstrated in the experiments (see Chapter 5). To achieve this, a self-supervised learning task is employed, which educates the model to formulate a representative and unique feature vector for an optimization instance – a process often termed feature-learning. Importantly, Deep-ELA is not restricted to single-objective optimization problems and can naturally be applied to multi-objective optimization problems. We will showcase Deep-ELA’s competitive edge against the traditional ELA but also feature-free approaches in existing studies (e.g., Seiler et al.,, 2022; Prager et al.,, 2022). We also applied our model to a bi-objective case analogous to Rook et al., (2022). See Chapter 5 for an outline of all the case studies.
To the best of our knowledge, only one alternative approach, termed DoE2Vec (van Stein et al.,, 2023), parallels our proposed method. However, DoE2Vec varies substantially from Deep-ELA: its model, rooted in a basic autoencoder design, solely focuses on the objective space $\mathcal{Y}$, completely disregarding the decision space $\mathcal{X}$. This reduced informational scope results in a markedly inferior performance compared to the approaches by Seiler et al., (2022). Additionally, the proposed method is not invariant to the order of objective values in $\mathcal{Y}$, potentially affecting its efficacy.
In terms of Deep-ELA’s architecture (refer to Figure 4), we predominantly adhered to the traditional transformer encoder blueprint as suggested by Vaswani et al., (2017). Contrary to the PCT design by Guo et al., (2021) and Seiler et al., (2022), we have embedded extra feed-forward modules to align more closely with the original transformer. Subsequently, we adopted pre-normalization as recommended by Nguyen and Salazar, (2019) and used Gated Linear Units (GLU; Dauphin et al.,, 2017) as the activation function between all feed-forward modules to reduce the training time. For the feature extractor (the last linear layer), we opted for the Tanh activation, mapping all activations to $[-1,1]$ and, hence, dispensing the need for feature normalization in later tasks.
The final model, Deep-ELA, can be formally expressed as
[…]
with ${c_{\text{D-ELA}}}$-dimensional feature space $\mathcal{F}_{\text{D-ELA}}\subseteq\mathbb{R}^{c_{\text{D-ELA}}}$. Analogous to ELA, Deep-ELA requires a sample of $(X,Y)\in(\mathcal{X},\mathcal{Y})$ as input. The resulting features lie within $[-1,1]$ and exhibit minimal correlation and redundancy. This contrasts with features from $\mathcal{F}_{\text{ELA}}$, which are un-normalized and often redundant (see Figure 2 for a comparison).
During the training phase, $X$ is drawn randomly uniformly within the bounded decision space $\mathcal{X}$. We favored uniform sampling over quasi-random sampling strategies to (1.) reduce additional computational overhead during the training procedure, (2.) reduce Deep-ELA’s dependence on a particular sampling method, and (3.) enhance the challenge of the self-supervised training task (which is described in more detail in Chapter 3.1.2). After pre-training, more advanced sampling methods, such as LHS and Sobol sampling, can be employed to improve coverage of the decision space.
Deep-ELA can accommodate any sample length of $(X,Y)$ but has a constraint regarding the sum of the search and objective space dimensions: $d+m\leq\nu$. Here, $\nu$ is an additional hyperparameter of Deep-ELA, termed the degree of dimensionality. It specifies the maximum combined dimensionality of $\mathcal{X}$ (denoted by $d$) and $\mathcal{Y}$ (denoted by $m$). Deep-ELA is versatile and can handle varying dimensionalities $d\in[1,\nu]$ and $m\in[1,\nu]$, provided their combined dimensionality does not surpass $\nu$. This flexibility extends to multi-objective optimization problems since the model can accept $m\geq 1$. It is worth noting that the model was trained on problem instances with $d\geq 2$ and $m\geq 1$.
3.1 Model Structure
[…]
Figure 4: Illustration of the chosen topology of the backbone model without the training heads. The model receives $(\mathcal{X},\mathcal{Y})$ as input and outputs $\mathcal{F}_{\text{D-ELA}}$ – and, optionally, the embedding of the tokens $\mathcal{T}_{\text{Final}}$ after the final LayerNorm. $\mathcal{T}_{\text{Final}}$ is only relevant for the contrastive loss during training and is ignored after training. The initial $k$NN embedding is used to capture the local information of all points from the input sample, and followed by a stride operator to optionally reduce the number of tokens without losing information. Next, the model consists of six Multi-Head Attention blocks, followed by a Feed-Forward block of two successive Linear layers each. The LayerNorm layers are after the shortcuts as proposed by Nguyen and Salazar, (2019). We chose GLU activations in the Feed Forward layers with a 4×4$\times$4 × larger number of hidden neurons. The last Linear + GLU layer projects the high-dimensional embeddings into lower dimensions. Afterward, the mean over all tokens is computed and normalized into $[-1,1]$ by a Tanh activation.
Our proposed Deep-ELA model is depicted in Figure 4. The workflow begins with the input $(X,Y)\in(\mathcal{X},\mathcal{Y})$, which undergoes a $k$-Nearest-Neighborhood ($k$NN) embedding as illustrated in Figure 6. This embedding methodology, originally introduced by Seiler et al., (2022), seeks to incorporate the local neighborhood of every $\mathbf{x}\in X$. Given that transformers excel at capturing global patterns, but may overlook local nuances, the $k$NN embedding ensures local information is not ignored. Post-embedding, the encoding undergoes a projection to the deep learning network’s hidden number of features, typically denoted as $d_{\text{model}}$ (refer to Section 3.1.1). Subsequently, every member of $\mathbf{x}\in X$ alongside its $k$ nearest neighbors is termed as tokens, aligning with the terminology conventionally associated with transformers.
Following this linear projection, the sequence of tokens is sequentially processed by six combined modules of Multi-Head Attention (MHA) and Feed-Forward (FF), concluding with a final Layer Normalization (LN; Ba et al.,, 2016). Through experimentation, we determined that six iterations strike an optimal balance between model intricacy and performance. The output post-LN layer manifests as a normalized $(n\times d_{\text{model}})$-dimensional tensor, with $n$ representing the number of tokens. To derive the features $\mathcal{F}_{\text{D-ELA}}$, the final embedding is condensed via a linear layer with GLU activation, referred to as feature-extractor, into a $\mathbb{R}^{n\times{c_{\text{D-ELA}}}}$ tensor. Ultimately, a mean pooling step across the token sequence yields ${c_{\text{D-ELA}}}$ features, which are constrained within $[-1,1]$ by a concluding Tanh activation. An illustration of the model’s topology and an overview of its associated hyperparameters are given in Figure 4 and Table 1, respectively.
This architecture termed the backbone model, is versatile for diverse downstream tasks. However, as the backbone essentially functions as a feature generator, it lacks trainability. Consequently, to facilitate training, we incorporate (1.) two feature-heads (distinguishing one as a so-called student and the other one as a teacher), and (2.) an additional teacher feature-extractor, as visualized in Figure 5. The student components are updated via gradient descent, while the teacher counterparts are updated via exponential moving average (EMA). Both are integral for the self-supervised loss computation. The training loss strategy revolves around providing distinct, augmented versions of the same objective instance to both the teacher and student. Here, the teacher generates target projections from which the student gleans insights. However, we refrained from using two backbone models (one for the student and one for the teacher) as this would drastically increase the total amount of model parameters. So, the backbone model is the same for both the student and teacher. More details about the loss function and the training routine are provided in Chapter 3.1.2.
$\mathcal{T}_{\text{final}}$$\mathcal{L}(\mathcal{P},\mathcal{\dot{P}})$$\mathcal{P}$$\mathcal{\dot{P}}$rom BB.Student’s HeadTeacher’s HeadLinear + GLUAvg. + TanhLinearGLU + DropoutLinearGLU + DropoutLinearBatchNormLinear + GLUAvg. + TanhLinearGLU + DropoutLinearGLU + DropoutLinearBatchNorm||||| |Update through Exponential Moving Average
Figure 5: The two training heads on top of the backbone model. Note that the first two layers before the student’s head are part of the backbone model (from BB.). Both heads are removed after training. The student’s head gets updated by gradient descent while the momentum head is an old version of the student’s head and gets updated through EMA. The design follows closely the idea of Chen et al., (2021).
Moreover, our approach employs the InfoNCE (Oord et al.,, 2018) loss function, specifically tailored for self-supervised learning. Notably, our model is strongly influenced by the principles of Momentum Contrast for Unsupervised Visual Representation Learning (MoCo; Chen et al.,, 2021) in its third iteration. Diverging from MoCo V3, we abstained from using a concluding MLP predictor, having observed its detrimental impact on performance and convergence in our context. Instead, our choice was a concluding Batch Normalization (BN; Ioffe and Szegedy,, 2015) with running mean and standard deviation but without the scale and shift parameters as we aim for $z$-standardization of the output projection. Even though the Batch Normalization layer did not notably elevate our performance, it amplified differences between dissimilar functions within the feature space, making distinctions more pronounced.
Given their pivotal roles in our model, the $k$NN embedding layer and the contrastive loss function will be described in more detail in the subsequent sections.
3.1.1 $k$NN Embedding
$\mathbf{x}_{1}$$\mathbf{x}_{4}$$\mathbf{x}_{2}$$\mathbf{x}_{5}$$\mathbf{x}_{3}$ind $k$NN of $\mathbf{x}_{i}\in X$ and create tokens $\mathbf{t}_{i}\in T$$\overbrace{\phantom{xxxxx}}^{\textbf{\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}Global Context}}$$\mathbf{t}_{1}=\big{(}{\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\mathbf{x}_{1},\mathbf{y}_{1}},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}(\mathbf{x}_{4}-\mathbf{x}_{1}),(\mathbf{y}_{4}-\mathbf{y}_{1}),(\mathbf{x}_{2}-\mathbf{x}_{1}),(\mathbf{y}_{2}-\mathbf{y}_{1})\ldots}\big{)}^{\text{T}}$$\mathbf{t}_{2}=\big{(}{\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\mathbf{x}_{2},\mathbf{y}_{2}},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}(\mathbf{x}_{1}-\mathbf{x}_{2}),(\mathbf{y}_{1}-\mathbf{y}_{2}),(\mathbf{x}_{4}-\mathbf{x}_{2}),(\mathbf{y}_{4}-\mathbf{y}_{2})\ldots}\big{)}^{\text{T}}$$\mathbf{t}_{3}=\big{(}{\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\mathbf{x}_{3},\mathbf{y}_{3}},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}(\mathbf{x}_{4}-\mathbf{x}_{3}),(\mathbf{y}_{4}-\mathbf{y}_{3}),(\mathbf{x}_{5}-\mathbf{x}_{3}),(\mathbf{y}_{5}-\mathbf{y}_{3})\ldots}\big{)}^{\text{T}}$$\mathbf{t}_{4}=\big{(}{\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\mathbf{x}_{4},\mathbf{y}_{4}},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}(\mathbf{x}_{1}-\mathbf{x}_{4}),(\mathbf{y}_{1}-\mathbf{y}_{4}),(\mathbf{x}_{2}-\mathbf{x}_{4}),(\mathbf{y}_{2}-\mathbf{y}_{4})\ldots}\big{)}^{\text{T}}$$\mathbf{t}_{5}=\big{(}{\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\mathbf{x}_{5},\mathbf{y}_{5}},\underbrace{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}(\mathbf{x}_{2}-\mathbf{x}_{5}),(\mathbf{y}_{2}-\mathbf{y}_{5}),(\mathbf{x}_{1}-\mathbf{x}_{5}),(\mathbf{y}_{1}-\mathbf{y}_{5})\ldots}}_{\textbf{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}Local Context}}\big{)}^{\text{T}}$
Figure 6: For every $\mathbf{x}_{i}\in X$ in the decision space, the $(k-1)$ nearest neighbors are identified and together with their respective objective values combined into a single vector. Next, the $(k-1)$ nearest neighbors are projected into the local neighborhood but centered by $\mathbf{x}_{i}$ and $\mathbf{y}_{i}$, i.e., subtracting them from all its $(k-1)$ nearest neighbors. We refer to $(\mathbf{x}_{i},\mathbf{y}_{i})$ as the global context and to all its $k-1$ neighbors as the local context.
Our Deep-ELA model, schematically represented in Figure 4, operates in two steps: input processing and embedding. Both are outlined in the following.
Input Processing:
The input sets $(X,Y)$ undergo individual $z$-standardization, both per set and per dimension. To extend their dimensionality, they are padded with zeros resulting in $X^{\prime},Y^{\prime}\in\mathbb{R}^{n\times\nu}$. They are then combined into a single set, $\dot{T}=X^{\prime}\,||\,Y^{\prime}$, where ||||| | indicates the element-wise concatenation operator. Following this procedure, $\dot{T}\subseteq\mathbb{R}^{n\times 2\nu}$ is structured such that its first $\nu$ dimensions are always decision values, while the subsequent $\nu$ dimensions represent objective values. Although this approach results in a matrix where at least half of the entries are zeros (given the dimensionality constraint $d+m\leq\nu$), it offers two advantages compared to using an additional indicator vector that discriminates between decision and objective values: (1.) all input variables remain real-valued, eliminating a mix of numeric and categorical values (that an identifier would introduce); (2.) adding an indicator vector of length $\nu$ would also produce an input of dimensionality $n\times 2\nu$.
Embedding:
For each $(\mathbf{x},\mathbf{y})\in(X,Y)$, its $k-1$ nearest neighbors are identified, projected to its local neighborhood, and then concatenated to $(\mathbf{x},\mathbf{y})$, resulting in:
$\mathbf{t}=\big{(}\mathbf{x},\mathbf{y},(\mathbf{x}_{1}-\mathbf{x}),(\mathbf{y}_{1}-\mathbf{y}),\ldots,(\mathbf{x}_{k-1}-\mathbf{x}),(\mathbf{y}_{k-1}-\mathbf{y})\big{)}^{\top}.$ (8)
The tokens $\mathbf{t}\in T$ form a vector with $T\subseteq\mathbb{R}^{n\times 2k\nu}$. Every element in $(X,Y)$ is termed as the global context due to its absolute position in the decision and objective spaces. Conversely, the $(k-1)$ nearest neighbors represent the local context, defined by their relative positions to each $(x,y)$. This distinction is pictorially elaborated in Figure 6. The embedding process can be formally outlined as:
$k\text{NN-Emb.}\colon(X,Y)\rightarrow\mathcal{T}\subseteq\mathbb{R}^{n\times 2k\nu}\text{.}$ (9)
Next, $\mathcal{T}$ undergoes a linear projection, followed by a token-wise applied GLU activation, mapping the points to a higher-dimensional space $\mathbb{R}^{n\times d_{\text{model}}}$ (see Figure 4).
3.1.2 Contrastive Loss
As previously highlighted, the model employs two distinct heads during training. The student head updates via gradient descent, while the teacher head employs exponential moving average (EMA) for updates. The outputs of these heads are termed as the online-projection $\mathcal{P}\subseteq\mathbb{R}^{8\cdot{c_{\text{D-ELA}}}}$ (from the student), and the target-projection $\dot{\mathcal{P}}\subseteq\mathbb{R}^{8\cdot{c_{\text{D-ELA}}}}$ (from the teacher). The eight-fold dimensionality of the projections increases the heads’ flexibility in aligning their representations. This methodology takes inspiration from Grill et al., (2020), who also considered higher dimensionalities for their heads.
The model’s training closely mirrors the principles of MoCo V3, including the adoption of the InfoNCE loss function, as introduced by Oord et al., (2018). It has shown state-of-the-art results in other feature-learning endeavors, as evidenced by Grill et al., (2020) and Chen et al., (2021). Given a batch of online-projections, $P=\big{\{}\mathbf{p}_{i}\in\mathcal{P}\mid\forall i\in\{1,\ldots,j\}\big{\}}$, and a batch of target-projections, $\dot{P}=\big{\{}\dot{\mathbf{p}}_{i}\in\dot{\mathcal{P}}\mid\forall i\in\{1,\ldots,j\}\big{\}}$, with $j$ being the batch-size, the overall goal of the InfoNCE is to fulfill the following equation:
$\sigma(P\dot{P}^{\top})\overset{!}{=}I\text{.}$
Here, $\sigma$ is the Softmax activation, and $I$ is the identity matrix. Broadly speaking, the loss function aims to maximize the covariance between an instance’s online- and target-projection and to minimize it between different instances. The InfoNCE loss function is formulated based on the cross-entropy:
$\mathcal{L}_{\text{InfoNCE}}(\mathcal{P},\dot{\mathcal{P}};\tau)=2\tau\cdot H\biggl{(}\sigma\bigg{(}\frac{\mathcal{P}\times\dot{\mathcal{P}}^{\text{T}}}{\tau}\bigg{)}\biggr{)}+D_{\text{KL}}\biggl{(}\sigma\bigg{(}\frac{\mathcal{P}\times\dot{\mathcal{P}}^{\text{T}}}{\tau}\bigg{)},I\biggr{)}\text{.}$ (10)
In this equation, $H$ and $D_{\text{KL}}$ denote the entropy and the Kullback-Leibler divergence, respectively. The cross-entropy approach provides pronounced gradients, even for vanishing activations, compared to the e.g. mean squared error (MSE). The loss function also depends on the critical hyperparameter $\tau$, which is the temperature of the Softmax activation $\sigma$. Usually, the value of $\tau$ is set in the range $(0,0.3]$. The smaller $\tau$ is, the higher the loss for hard negative samples, i.e., distinct instances that yield similar projections and, thus, are hard for the model to differentiate. On the contrary, for larger values of $\tau$, the loss is more evenly spread over hard and easy negative samples. Through initial testing, we found that $\tau=0.05$ works well in our scenario. To increase the likelihood of the appearance of hard negative samples within a single batch, the batch size is supposed to be as large as possible, which is why we aimed for very large training batches.
During training, the model receives two augmented pairs of $(X_{1},Y_{1}),(X_{2},Y_{2})$ for every problem instance $Z$ in a batch of $j$ problem instances. For every batch, the student and also the teacher process both augmentations of every instance, yielding four distinct projections: $P_{1},P_{2}\in\mathcal{P}$ (from the student) and $\dot{P}_{1},\dot{P}_{2}\in\dot{\mathcal{P}}$ (from the teacher). The overall loss is computed as:
$\mathcal{L}=\frac{1}{2}\big{(}\mathcal{L}_{\text{InfoNCE}}(P_{1},\dot{P}_{2};\tau)+\mathcal{L}_{\text{InfoNCE}}(\dot{P}_{1},P_{2};\tau)\big{)}\text{.}$ (11)
The deliberate pairing of $(P_{1},\dot{P}_{2})$ and $(\dot{P}_{1},P_{2})$ aims to maximize the loss across both augmented versions. For augmentation, we adopted rotations and inversions of decision variables and randomized the sequence of decision and objective variables. As the augmentations do not alter the underlying optimization problem, the predicted features should be invariant to these changes. Furthermore, the $z$-standardization in the embedding layer ensures the model remains invariant to scale and shift modifications.
4 Datasets
In total, we considered four sets of optimization problem suites for training: single-objective problems from the Black-Box Optimization Benchmarking Suite (BBOB; Hansen et al.,, 2009), multi-objective instances from the Biobjective-BBOB Suite (Bi-BBOB; Brockhoff et al.,, 2022), selected instances from the R-package smoof [^1] https://github.com/jakobbossek/smoof (Single- and Multi-Objective Optimization Test Functions), and an adapted version of the random function generator as introduced by Tian et al., (2020).
4.1 Black-Box Optimization Problems
Most validation and testing optimization problems come from BBOB, provided by the COmparing Continuous Optimisers (COCO; Hansen et al.,, 2019) platform. COCO offers a diverse range of continuous optimization problem suites, spanning from single- and multi-objective to noisy and mixed-integer problems. This study primarily engaged with the 24 noiseless, single-objective functions from the (classical) BBOB suite. For each of these functions, 1 00010001\,0001 000 unique instances were produced, of which the first 500 served for testing and the latter for validation purposes (see Chapter 5). Function dimensionalities were chosen from $d\in\{2,3,5,10\}$. Although the suite includes functions with $d\in\{20,40\}$, our model’s current dimensional constraints prevent their inclusion.
For the multi-objective case study, we analyzed the same multi-objective problems that Rook et al., (2022) examined. These bi-objective ($m=2$) functions have a decision space dimensionality of $d=2$. Specifically, we considered the Bi-BBOB functions $f_{46},f_{47},$ and $f_{50}$, along with the ZDT, DTLZ, and MMF function groups (excluding ZDT5 and MMF13) from the smoof package. Further details are given in Chapter 5.4.
4.2 Random Optimization Problems
Refer to caption (a) Dispersion
Refer to caption (b) Meta
Refer to caption (c) $Y$-Distribution
Figure 7: Distributions of feature values for three exemplary sets of ELA features, differentiated by three types of test problems – the original instance generator by Tian et al., (2020) (blue), the BBOB suite (orange), and our adapted generator (green). For all problems, we used dimensions $2,3,5,10$. In the case of BBOB, we considered all 24 noiseless functions with instances $1$ to $20$. For the random generators, we used the same number of instances (in total).
Tian et al., (2020) introduced a method capable of generating an arbitrary number of optimization problems in milliseconds. This approach hinges on a random tree-like structure constituted of various operators such as mean, sum, exp, log, among many others. This generator operates on three input parameters: the number of dimensions of the decision space ($d$), and the lower and upper bounds of the number of operators. Once initiated, the generator fabricates an optimization problem for a $d$-dimensional set of decision variables $\mathcal{X}$ by randomly assorting a number of operators bounded by the provided lower and upper bounds. The result is a randomly constructed optimization problem that yields the objective values $Y$. We manually calibrated the generator’s hyperparameters to craft optimization problems that are similar to BBOB. Additionally, we incorporated extra operators to amplify the linear and squared dependencies among decision and objective variables, noting that the original generator somewhat diverged from the BBOB suite in these aspects. Furthermore, we tweaked the skewness and kurtosis of the objective variables for closer alignment with BBOB (refer to Figure 7).
However, given the inherent randomness of the generator, not all produced optimization problems are applicable for training. We thus dismiss any instance that does not satisfy the following three conditions:
  1. 1.
    The output $Y$ must be a set of scalars.
  2. 2.
    The objective values must have a standard deviation of at least $0.1$
  3. 3.
    The objective values must lie within [10 000 000,+10 000 000]1000000010000000[-10\,000\,000,\;+10\,000\,000][ - 10 000 000 , + 10 000 000 ].
To generate multi-objective instances, we randomly designed $m$ single-objective functions and combined them into a multi-objective instance. While this methodology proved satisfactory for our purposes, the combined set of random single-objectives may not fully encapsulate the inherent traits of multi-objective problems. Notably, such problems often harbor interacting objectives, potentially exhibiting mutual dependencies. However, randomly creating multi-objective instances with such interdependencies proved to be much more challenging. We thus postponed this endeavor for future work, and hope that randomness will generate objectives with mutual dependencies. We built upon the Python implementation [^2] https://www.github.com/Basvanstein/doe2vec/tree/main/src/modulesRandFunc by van Stein et al., (2023).
5 Experiments
In the following sections, we describe and analyze the pre-training of our Deep-ELA models. Subsequently, we examine the results of three case studies. The first two, High-Level Property Prediction (Seiler et al.,, 2022) and Single-Objective Automated Algorithm Selection (Prager et al.,, 2022), have been conducted in previous studies. Please note that we directly took the results from previous work of Seiler et al., (2022); Prager et al., (2022); van Stein et al., (2023). The third case study, i.e., the Multi-Objective Automated Algorithm Selection, is adapted from Rook et al., (2022) and introduces a newly created dataset.
5.1 Pretraining
Table 1: Hyperparameter settings of our models. In total, we trained four final models, which were configured systematically based on a pre-study. The number of randomly generated training instances is given per epoch, so, every model was trained on a total of 250 000 000 randomly generated single- and multi-objective functions.
Parameter Medium Model Large Model
Degree of Dim. $25d$ $50d$ $25d$ $50d$
$\nu$ 6 6 12 12
$k$ 8 8 16 16
Num. MHA 6 6 6 6
Num. Heads 4 4 8 8
$\text{d}_{\text{model}}$ 192 192 384 384
Epochs 250 250 250 250
Train Instances 1 000 000 1 000 000 1 000 000 1 000 000
Batch Size 1 264 632 1 264 632
Acc. Grad 1 2 1 2
$\tau$ 0.05 0.05 0.05 0.05
Stride 1 1 2 2
EMA Momentum 0.01 0.01 0.01 0.01
BN Momentum 0.1 0.1 0.1 0.1
Device 3×3$\times$3 × NVidia Q. RTX 6 000 3×3$\times$3 × NVidia RTX A 6 000
GPU Memory 3×3$\times$3 × 24GB 3×3$\times$3 × 48GB
Precision Float16 (mixed) BFloat16 (mixed)
Params. BB. 2 263 296 9 189 888
Params. Total 2 355 456 9 558 528
Before training our final models, we conducted several preliminary studies to determine the optimal hyperparameters and model topologies for our use cases. We experimented with various $\tau$ values, momentum factors, numbers of heads, and layers. Eventually, we systematically selected the hyperparameters presented in Table 1.
Given the availability of two high-performance computing (HPC) servers – one with three NVidia Quadro RTX 6 000 GPUs and the other with three NVidia RTX A 6 000 GPUs – we decided to train the larger models on the RTX A 6 000 GPUs, which have 48GB of GPU memory, compared to the Quadro’s 24GB. The smaller models were comfortably trained on the 24GB GPUs. We maximized the batch size, retaining a small buffer, to include as many samples as possible within each batch as the used loss function benefits from large batches. Each model was trained across three GPUs, handling three batches concurrently. After the forward pass, the three batches were merged into one joint batch for loss computation, effectively tripling the batch size and increasing the likelihood of the appearance of hard negative samples.
We chose to train both a medium- and a large-sized model. The medium model can manage a total dimensionality of six, while the large model can accommodate up to twelve total dimensions. The large model has twice the width of the medium model and produces twice the number of features (48 vs. 24). Due to the squared growth of transformer complexity with the number of tokens, we introduced a stride factor. A stride of two implies that every second token is omitted after the $k$NN-embedding, as depicted in Figure 4. Consequently, the large model’s complexity is reduced to a fourth, compared to a model with a stride of one, maintaining a larger batch size during the training of the large model. We then trained each model-variant with sample sizes of $25d$ and $50d$, resulting in four final models. For the $50d$ models, we reduced the batch size by half but compensated with gradient accumulation over two iterations. However, InfoNCE does not benefit much from gradient accumulation since its efficiency relies on a substantial number of samples. Still, we opted for gradient accumulation to maintain a consistent number of update steps and training instances, as seen with the $25d$ models.
For training, we employed PyTorch-Lightning [^3] https://lightning.ai/ in tandem with PyTorch [^4] https://pytorch.org/. Post-training, we removed the two heads necessary during the training phase, resulting in a versatile backbone model suitable for various downstream tasks. Subsequently, we incorporated classical machine learners atop this backbone. No further fine-tuning of the backbone model was undertaken or deemed necessary. However, it is possible to fine-tune the backbone model for different tasks, although this demands a deep understanding of deep learning techniques. To emulate a user proficient in general machine learning but not deeply versed in deep learning specifics, we relied exclusively on Scikit-Learn and refrained from fine-tuning the backbone model. Further, we did not apply feature selection to demonstrate that the instance features generated by the Deep-ELA approach contain features of high relevance and little redundancy.
5.2 High-Level Property Prediction
Table 2: Assignment of the 24 single-objective noiseless BBOB functions into three frequently used High-Level Properties (HLP), as outlined and proposed in Mersmann et al., (2010) and Kerschke et al., (2015).
BBOB Function Multim. Global Str. Funnel
1: Sphere none none yes
2: Ellipsoidal separable none none yes
3: Rastrigin separable high strong yes
4: Büche-Rastrigin high strong yes
5: Linear Slope none none yes
6: Attractive Sector none none yes
7: Step Ellipsoidal none none yes
8: Rosenbrock low none yes
9: Rosenbrock rotated low none yes
10: Ellipsoidal high conditioned none none yes
11: Discus none none yes
12: Bent Cigar none none yes
13: Sharp Ridge none none yes
14: Different Powers none none yes
15: Rastrigin multimodal high strong yes
16: Weierstrass high med. none
17: Schaffer F7 high med. yes
18: Schaffer F7 moderately ill-cond. high med. yes
19: Griewank-Rosenbrock high strong yes
20: Schwefel med. deceptive yes
21: Gallagher 101 Peaks med. none none
22: Gallagher 21 Peaks low none none
23: Katsuura high none none
24: Lunacek bi-Rastrigin high weak yes
The term High-Level Properties (HLP) refers to a collection of structural attributes describing the fitness landscape of the examined single-objective optimization. These properties enable the characterization of the problem – e.g., into low, medium, and high multimodality – facilitating the identification of (dis-)similar problems. This holds especially true for high-dimensional problems, which cannot be visually represented or analyzed. Moreover, these properties are pivotal for algorithm selection, algorithm configuration, and for crafting algorithms tailored to specific objectives. Of the eight properties detailed in Mersmann et al., (2011) and Kerschke et al., (2015), the following three exert the most pronounced influence on optimization problem complexity:
  1. 1.
    Multimodality: The degree of multimodality aggregates the number of local optima into ‘low’, ‘medium’, and ‘high’ multimodality, as well as ‘none’ in case of unimodal problems.
  2. 2.
    Global Structure: Chronicles the structural link between local and global optima.
  3. 3.
    Funnel: Signifies the presence of a funnel-like layout of the local and global optima.
The assignment of the 24 BBOB functions into the three HLPs is given in Table 2 for convenience. For predicting the HLPs, we adhered to the setup prescribed in Seiler et al., (2022). This entailed evaluating the 24 BBOB functions, adopting instances $\{1,\ldots,100\}$ for training and $\{125,\ldots,150\}$ for testing, with $d\in\{2,3,5,10\}$. We juxtaposed our findings against those from Seiler et al., (2022). Additionally, we consider the results reported by van Stein et al., (2023) in the HLP study of their Doe2Vec method. For prediction purposes, we utilized Random Forests (RF) and Support Vector Machines (SVM) to forecast the three HLPs based on features generated by the pre-trained backbone model. The obtained (and collected) results are listed in Table 3.
Table 3: F1-Score with macro aggregation (F1-Scores a computed individually per class and then mean-aggregated over all classes) of predicting High-Level Properties of all 24 BBOB functions and instances $126$ to $150$. The results marked with a $*{}^{*}$ are directly taken from Seiler et al., (2022) while the ones marked with ${}^{**}$ are taken from van Stein et al., (2023). The last eight columns list the performances obtained by our proposed Deep-ELA models. van Stein et al., (2023) did not provide results for $d=3$ and aggregated over all dimensions (all). Our medium models cannot handle data with more than six total dimensions. Therefore, there are no results for $d=10$ for the medium models.
[…]
High-Level Dim. ($50d$) p100 p500 ($50d$) ($50d$) RF SVM RF SVM RF SVM RF SVM
Multi- 2 0.997 0.991 0.997 0.849 0.856 0.854 0.845 0.895 0.904 0.896 0.905 0.938 0.939
modality 3 0.997 0.988 0.994 -/- -/- 0.869 0.877 0.946 0.953 0.958 0.950 0.974 0.973
5 0.999 0.991 0.999 0.903 0.889 0.922 0.942 0.945 0.953 0.959 0.966 0.961 0.966
10 1.000 0.974 0.991 0.813 0.838 0.920 0.915 0.941 0.949 -/- -/- -/- -/-
all 0.998 0.986 0.995 -/- -/- 0.893 0.896 0.932 0.940 0.938 0.940 0.958 0.959
Global- 2 0.997 0.991 0.998 0.904 0.889 0.857 0.849 0.921 0.941 0.935 0.946 0.953 0.956
Structure 3 0.996 0.986 0.994 -/- -/- 0.890 0.909 0.954 0.968 0.960 0.961 0.976 0.977
5 0.998 0.978 0.995 0.828 0.793 0.926 0.936 0.948 0.940 0.954 0.957 0.957 0.957
10 0.999 0.963 0.984 0.737 0.745 0.898 0.887 0.933 0.953 -/- -/- -/- -/-
all 0.998 0.980 0.993 -/- -/- 0.894 0.895 0.939 0.950 0.950 0.955 0.962 0.963
Funnel- 2 0.998 0.999 1.000 0.974 0.978 0.986 0.980 0.994 0.994 0.988 0.991 0.994 0.991
Structure 3 1.000 1.000 1.000 -/- -/- 1.000 0.994 1.000 1.000 1.000 0.997 1.000 1.000
5 1.000 1.000 1.000 1.000 0.991 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
10 1.000 0.999 1.000 0.993 0.993 1.000 1.000 1.000 1.000 -/- -/- -/- -/-
all 1.000 1.000 1.000 -/- -/- 0.996 0.994 0.999 0.999 0.996 0.996 0.998 0.997
As summarized in Table 3, the performances of our models lag behind those of Seiler et al., (2022) but are superior to the ones listed in van Stein et al., (2023). This was somewhat anticipated, given that Seiler et al., (2022) explicitly trained on BBOB instances for HLP prediction. In contrast, both van Stein et al., (2023) and our approach were trained by self-supervised learning on autonomously generated instances. Notably, our medium-scale models often outperformed their larger counterparts. Two potential reasons underpin this trend. Firstly, the large models, having been utilized with a stride of two, bypass half of the accessible data, potentially decreasing their efficacy. The second rationale suggests that while the larger models underwent an identical number of training iterations, they were exposed to a broader range of dimensions, possibly hindering their ability to generalize equally well to specific dimensions. Lastly, our models with $50d$ consistently outperformed the $25d$ models, echoing the findings of Seiler et al., (2022), who observed superior performance from transformers with a sample size of 500 versus those with a size of 100. In terms of the considered machine learning models, our experiments did not discern any substantial performance discrepancy between the RF and SVM results; both exhibited roughly comparable efficiency.
5.3 Single-Objective Automated Algorithm Selection
In our subsequent case study, we both replicated and expanded upon the experiments presented by Prager et al., (2022). Integral to this study was the data originally compiled and introduced by Kerschke et al., (2015), encompassing benchmark results of twelve competitive, complementary algorithms – a comprehensive list of these algorithms can be found in Kerschke et al., (2015) and Prager et al., (2022). The benchmark performances were obtained from COCO, a repository archiving the outcomes of optimizer runs from a myriad of competitions hosted on the BBOB suites. In the methodology laid out by Kerschke et al., (2015), a threshold of $0.01$ was deemed acceptable, i.e., an algorithm that procures a solution within $0.01$ of the actual optimum (in objective space) is categorized as successful. An algorithm’s efficacy is gauged through the relative Expected Running Time (relERT) metric (Auger and Hansen,, 2005). The relERT provides a relative variant of the ERT itself, by juxtaposing the ERT of a given algorithm against that of the Virtual Best Solver (VBS). Hence, a relERT of $r$ indicates that the examined algorithm needs $r$ times as many function evaluations compared to the best algorithm from the considered portfolio. Note that when predicting the top-performing algorithm, the size of the initial sample must be factored in as a cost, which is first added to the algorithm’s ERT prior to deriving the relERT. The sample is essentially needed for the model to scrutinize a given problem instance (in feature-free algorithm selection) and predominantly for feature computation (in feature-based algorithm selection).
Table 4: Comparison of the relative ERT (relERT) values for the algorithm selection study on instances $1$ to $5$ of the 24 single-objective BBOB functions. The first five columns (marked with $*{}^{*}$) are directly taken from Prager et al., (2022), while the last eight show the performance of our proposed Deep-ELA approach. Our medium models cannot handle data with more than six dimensions. Therefore, there are no results for $d=10$ for the medium models.
[…]
D F. Group $*{}^{*}$ RF MLP p100 p500 $k$NN RF $k$NN RF $k$NN RF $k$NN RF
2 1 3.71 10.41 10.59 23.95 61.15 9.24 10.30 14.26 13.87 17.57 7.57 16.38 14.77
2 5.80 8.51 3.72 3.54 11.04 2.65 2.87 3.49 4.25 2.69 3.37 5.70 4.08
3 6.29 1473.16 4.72 4.02 16.11 3.52 3.12 5.32 4.80 3.91 4.40 5.33 4.65
4 25.34 3.89 9.25 8.90 12.21 6.45 9.76 5.64 6.76 5.73 5.99 3.48 4.29
5 44.95 148.39 3.32 4.33 6.18 3.69 5.06 3.81 7.77 3.79 8.08 3.45 14.58
all 17.69 342.22 6.43 9.17 21.77 5.21 6.36 6.63 7.62 6.91 5.99 6.92 8.66
3 1 356.10 1480.68 11.87 13.32 39.24 19.94 66.76 15.29 15.54 11.72 53.23 15.32 16.13
2 4.46 8.33 3.50 2.85 6.65 2.73 3.02 3.74 3.75 2.67 2.87 3.42 3.62
3 4.98 7.07 3.82 2.72 9.58 2.69 3.48 3.72 4.81 2.59 3.53 3.85 4.02
4 2.63 441.96 5.06 11.49 11.87 5.15 4.63 5.20 5.00 5.36 4.83 5.12 4.20
5 66.81 1.22 2.54 2.46 3.04 30.75 12.02 4.21 25.80 2.70 4.99 7.24 6.85
all 90.43 403.67 5.44 6.72 14.39 12.65 18.61 6.54 11.28 5.10 14.35 7.14 7.10
5 1 11.99 14.14 11.97 16.27 33.39 17.70 17.31 22.88 22.81 17.50 17.85 24.18 24.01
2 3.90 369.26 2.62 4.25 5.66 2.40 4.27 3.03 3.89 2.48 2.45 3.59 3.39
3 4.21 150.44 3.97 5.21 9.23 3.70 4.87 4.63 6.29 3.76 3.69 4.33 4.58
4 4.29 1470.28 6.81 4.33 4.47 1.87 4.07 1.90 4.24 3.95 3.51 2.25 2.44
5 7.67 1.13 1.83 7.72 7.93 1.08 4.73 1.15 3.71 1.72 1.43 1.57 1.96
all 6.52 402.38 5.56 7.69 12.41 5.47 7.17 6.87 8.37 6.02 5.93 7.33 7.44
10 1 2.74 14.64 15.27 5.46 16.34 9.54 9.53 16.45 16.28 -/- -/- -/- -/-
2 2.16 1.62 1.76 2.27 2.71 2.40 2.43 2.64 2.71 -/- -/- -/- -/-
3 2.76 2.87 4.35 3.10 4.48 3.54 3.62 4.52 4.15 -/- -/- -/- -/-
4 2.02 442.01 1.96 2.03 2.09 2.06 2.05 1.91 2.09 -/- -/- -/- -/-
5 23.64 148.01 3.25 23.66 23.74 1.73 11.33 1.80 12.09 -/- -/- -/- -/-
all 6.85 126.84 5.46 7.51 10.17 3.91 5.93 5.58 7.66 -/- -/- -/- -/-
all 1 93.63 379.97 12.43 14.75 37.53 14.10 25.97 17.22 17.13 15.60 26.22 18.62 18.30
2 4.08 96.93 2.90 3.23 6.52 2.54 3.15 3.23 3.65 2.62 2.90 4.24 3.69
3 4.56 408.38 4.21 3.76 9.85 3.36 3.77 4.55 5.01 3.42 3.88 4.50 4.42
4 8.57 589.54 5.77 6.69 7.66 3.88 5.13 3.66 4.52 5.01 4.78 3.62 3.64
5 35.77 74.69 2.74 9.54 10.22 9.31 8.29 2.74 12.34 2.73 4.83 4.09 7.80
all 30.37 318.78 5.72 7.78 14.68 6.81 9.52 6.41 8.73 6.01 8.76 7.13 7.73
Our experimental setup closely mirrored that of Prager et al., (2022), employing the 24 noiseless BBOB functions, instances $\{1,2,3,4,5\}$ with ten repetitions each, and $d\in\{2,3,5,10\}$. In line with Prager et al., (2022), we executed five-fold cross-validation across the five instances for each function. Further, Prager et al., (2022) employed a cost-sensitive loss function, but it was exclusively compatible with machine learners amenable to gradient descent training. This discrepancy is evident when comparing the RF and MLP models, both trained on ELA features. While the RF model significantly underperformed in relation to the Single Best Solver (SBS), the MLP exhibited state-of-the-art results. Our research, however, sidestepped the loss function delineated by Prager et al., (2022), as we restricted ourselves to the sklearn framework, avoiding other platforms. In contrast, Prager et al., (2022) utilized PyTorch complemented by a bespoke loss function and training protocol. sklearn, however, does not provide the functionality to implement a custom loss function or to make use of an instance-based, cost-sensitive training routine.
Diverging from the insights of our initial case study, this research unveiled a noticeable enhancement in our proposed Deep-ELA methodology compared to the results of Prager et al., (2022). All our models consistently outperformed the feature-free strategies delineated in Prager et al., (2022). In certain scenarios, our Large-$25d$ and Medium-$50d$ models even surpassed the MLP-ELA model. A consistent observation from Prager et al., (2022) was the superior performance of the Medium-$25d$ over the Medium-$50d$ model. We postulate that the higher costs associated with acquiring supplementary information outweigh the benefits derived from this extra data. However, this paradigm does not apply to the large models, with the Large-$50d$ model being superior to the Large-$25d$ model. This trend suggests that larger models may be necessary to extract additional value from the added data of the larger sample size. Nevertheless, the disparity between the results is nuanced. Our best-performing Deep-ELA model was the Medium-$25d$ model, lending credence to the hypothesis that employing a stride of two may hamper performance. Notably, the $k$NN classifier substantially outperformed the RF classifier. In this context, we opted for a $k$ value of 69, the sole hyperparameter we carefully fine-tuned manually.
Our investigations into single-objective AAS unveiled a pivotal revelation: Deep-ELA operates as envisaged, particularly excelling with limited datasets compared to feature-free AAS, where the propensity for overfitting makes training extensive deep-learning models a challenge.
5.4 Multi-Objective Automated Algorithm Selection
Our experimental design is akin to that of Rook et al., (2022), albeit without the algorithm configuration segment. Instead, we opted for the seven algorithms (as detailed in Rook et al.,, 2022) in their default settings and employed an AAS approach to select the optimal algorithm. Three of these seven algorithms – NSGA-II (Deb et al.,, 2002), SMS-EMOA (Beume et al.,, 2007), and MOEA/D (Zhang and Li,, 2007) – represent classical Evolutionary Multiobjective Optimization Algorithms (EMOAs), while the other four are Omni-Optimizer (Deb and Tiwari,, 2005), MOLE (Schäpermeier,, 2022), MOGSA (Grimme et al.,, 2019), and HIGA-MO (Wang et al.,, 2017). Performance was gauged using the Hypervolume (HV) of the approximated Pareto front found within a budget of 20 0002000020\,00020 000 function evaluations. For most test instances, the reference point for the HV was pre-specified. For problems without predetermined reference points (for HV computation), we derived the necessary points from the least favorable solution of all algorithms for that specific instance. To maintain HV value consistency across distinct instances, we normalized the HV values, basing them on the highest HV identified from additional runs of all algorithms with a more substantial budget of 100 000100000100\,000100 000 function evaluations. Subsequently, the resulting normalized HV values were contrasted against the SBS-VBS gap – and we coined this metric relative HV (relHV). In this context, a value close to zero resembles SBS performance, while a value of one relates to VBS performance. Negative values, in turn, denote performances that are inferior to the SBS. Formally, the metric is defined as
$\text{relHV}(I,A)$ $=\frac{\text{HV}(I,A)-\text{HV}_{\text{SBS}}+10^{-8}}{\text{HV}_{\text{VBS}}-\text{HV}_{\text{SBS}}+10^{-8}}$ (12)
where $10^{-8}$ is a tiny value that prevents division by zero. This could happen if the SBS and VBS are identical. By also including the term $10^{-8}$ in the numerator ensures that the relHV is one, in case the selector predicts the VBS.
Table 5: The mean relative HV (relHV) of our proposed Deep-ELA approach on the multi-objective AAS study. relHV values are calculated per test instance and mean-aggregated first over instances of the same function and then over functions of the same group. A value of one (zero) indicates performances comparable to the VBS (SBS). Negative values indicate a performance worse than the SBS.
[…]
Group $k$NN RF $k$NN RF $k$NN RF $k$NN RF
BiBBOB 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
DTL 0.914 0.457 1.000 0.651 1.000 0.651 1.000 0.822
MMF 0.892 0.463 0.931 0.609 0.929 0.764 1.000 0.933
ZDT -0.079 -0.087 0.041 0.209 0.688 0.185 0.616 0.745
all 0.682 0.458 0.743 0.617 0.904 0.650 0.904 0.875
The dataset comprised ZDT, DTLZ, and MMF test instances, excluding ZDT5 and MMF13. Additionally, instances $f_{46}$, $f_{47}$, and $f_{50}$ from the Bi-BBOB were integrated. This resulted in a total of 33 instances – with a two-dimensional decision ($d=2$) and objective ($m=2$) space each. We executed 20 repetitions per instance, of which the first 15 were used for training and the remaining five ($16-20$) for testing.
The results of our exploration are listed in Table 5. Due to the lack of other AAS studies on multi-objective optimization, we cannot compare Deep-ELA to any other studies from the literature. A potential reason for this may be the somewhat limited applicability of ELA on multi-objective optimization problems. It is pertinent to note that the crux of this study revolves around maximizing the relHV. Given the absence of any costs to factor in for this case study – as we measure the performance based on the found hypervolume instead of the number of function evaluations – achieving performances on the VBS level is viable. Our findings reveal that the medium models eclipse the larger ones in performance, and the $50d$ models have a performance edge over the $25d$ models. This observation aligns with our expectations, considering that the $50d$ models harness more information than their $25d$ counterparts. We posit that the larger models are potentially hampered by the stride of two. Specifically, the large $25d$ model registered a negative relVH value for the ZDT instance group, suggesting a performance that lags behind the SBS. The $k$NN classifiers showcased a discernibly superior performance in comparison to the RF classifiers. In this context, we opted for a smaller $k=15$. Both, the Medium $25d$ and Medium $50d$ models rendered comparable outcomes when leveraging a $k$NN classifier. However, when an RF classifier was in play, the Medium $50d$ model emerged as the top performer. In conclusion, the performances achieved by our Deep-ELA approach in this study are often very good – and, in many cases, even perfect.
6 Discussion & Conclusion
A salient outcome of this research is the demonstrated efficacy of the Deep-ELA methodology. It can either be used out-of-the-box for analyzing single- and multi-objective continuous optimization problems, as demonstrated, or fine-tuned to various tasks on algorithm behavior and problem understanding. In comparison to the strategies delineated by Prager et al., (2022), Deep-ELA exhibited significant performance enhancements, particularly in its adept handling of limited datasets. Feature-free deep-learning models, like those utilized by Prager et al., (2022), often contend with overfitting challenges when trained on limited datasets. In contrast, Deep-ELA’s design, having been trained on millions of randomly generated optimization instances, offers a remedy to such constraints. As a result, the backbone model could be integrated seamlessly into existing and novel downstream tasks without necessitating additional training or fine-tuning on the dataset from Prager et al., (2022). Furthermore, the derived instance features could be employed without the need for feature selection or normalization, as these features are less correlated than the commonly used ELA features, and, in addition, are all located within $[-1,1]$.
Yet, alongside these advantages, the investigations revealed certain intricacies warranting further examination. Notably, medium-sized models demonstrated superior performance compared to their larger counterparts. Concurrently, a trend emerged indicating that models with a sample size of $50d$ performed better than those with $25d$. While the latter trend is intuitive – a sample size of $50d$ inherently offers more information for the model to process – the former raises questions regarding the current iteration of the Deep-ELA design. The use of a stride of two, particularly in larger models, was pinpointed as a possible bottleneck inhibiting peak performance. This observation offers a compelling avenue for further research. Even though the chosen stride of size two was indispensable for accommodating large training batches, future studies could explore more refined pooling techniques to diminish sequence sizes without significant data loss. Another potential explanation for the diminished performance of larger models may lie in their training on a more varied set of decision space dimensions. As a future direction, researchers may delve into whether extended training durations would amplify results or if other underlying factors are at play.
Despite the contained scope of the results, the Deep-ELA methodology made a noteworthy foray into the multi-objective AAS domain. The capability of Deep-ELA to seamlessly transition to multi-objective problem instances stands out as a significant achievement. This is particularly notable considering that traditional ELA features are innately tailored for single-objective problems. Moving forward, there is ample opportunity to devise more intricate studies where the Deep-ELA approach can be further tested and refined. At last, configuring the hyperparameters of the Deep-ELA framework in an automated manner, e.g., using efficient algorithm configurators such as irace (López-Ibáñez et al.,, 2016) or SMAC (Lindauer et al.,, 2022), also provides a very promising direction for future research.